OpenStudy (anonymous):

Let A be the area of a circle with radius r. If \displaystyle \frac { dr }{ dt } = 4, find \displaystyle \frac { dA }{ dt } when r = 4

OpenStudy (anonymous):

\[8\Pi r\]

OpenStudy (anonymous):

\[A=\pi r^2\] \[A'=2\pi r r'\]

OpenStudy (anonymous):

\(r=4, r'=4\) you get \[2\pi\times 4\times 4 \]

OpenStudy (anonymous):

Thanks